# A Survey of Reinforcement Learning For Economics

### 23 Mar 2026 v4

@@ 8.3.2 Fitted Policy Iteration for Combinatorial Auctions
The mechanism’s policy maps from state to a price vector over available items. The key technical contribution is making the auction clearing differentiable. In a standard auction, the allocation is an argmax over bids (non-differentiable), and payments depend discontinuously on the allocation. Ravindranath et al. (2024) replace the hard bundle selection with a softmax relaxation,166166166The softmax relaxation replaces the hard allocation $\mathop{\it argmax}_{i}b_{i}$ with a soft allocation $\exp(b_{i}/\tau)/\sum_{j}\exp(b_{j}/\tau)$, which is differentiable and approaches the hard allocation as $\tau\to 0$. enabling analytical gradient computation through the mechanism. The actor loss is the negative expected revenue, and gradients flow directly through the softmax-relaxed allocation and payment rules. The paper explicitly avoids REINFORCE-style estimators, noting that analytical gradients overcome the sample inefficiency and high variance of score-function methods.
@@
Experiments on settings with additive and combinatorial valuations show that the learned mechanisms achieve up to 13% higher revenue than item-wise Myerson optimal auctions, with the largest gains in combinatorial settings where bundle complementarities make item-wise pricing suboptimal. Standard RL baselines (PPO, SAC) are also outperformed, confirming the advantage of exploiting differentiable structure.167167167DDPG was also tested but found to be unstable. DQN is not applicable to continuous price-setting.
@@ 8.4 Macroeconomic Models
Reinforcement learning and deep learning are increasingly used to solve high-dimensional macroeconomic models where grid-based dynamic programming is infeasible. Heterogeneous agent models, in which a distribution of agents with different wealth levels, productivities, or beliefs interact through markets, generate state spaces that grow with the number of agent types and asset positions. Maliar et al. (2021) demonstrate that deep neural networks can approximate policy and value functions in dynamic economic models, achieving accuracy comparable to established projection methods while scaling to problems with dozens of state variables. Fernández-Villaverde et al. (2023) solve a model with financial frictions and an endogenous wealth distribution using deep learning, obtaining global solutions to a problem where perturbation methods fail due to strong nonlinearities. Fernández-Villaverde et al. (2024) provide a systematic treatment of deep learning methods for high-dimensional dynamic programming problems in economics, covering both single-agent and equilibrium settings. Atashbar and Shi (2023) apply deep deterministic policy gradient (DDPG)168168168DDPG extends Q-learning to continuous action spaces by learning a deterministic policy network alongside a Q-function critic. to a real business cycle model, demonstrating that model-free RL can recover near-optimal consumption and investment policies without deriving optimality conditions such as the Euler equation. Moll (2025) argues that rational expectations equilibria in heterogeneous agent models are computationally intractable and proposes reinforcement learning as a more tractable alternative for modeling how agents form beliefs and make decisions; see also the textbook treatment in Zhao (2025). This is a rapidly growing area; readers seeking a comprehensive methodological treatment are directed to the cited papers and the references therein.
@@ 9.1.1 The Stochastic Game Framework
An $n$-player stochastic game $\Gamma=(n,\mathcal{S},\mathcal{A}_{1},\ldots,\mathcal{A}_{n},P,r_{1},\ldots,r_{n},\gamma)$ consists of a finite state space $\mathcal{S}$; finite action sets $\mathcal{A}_{i}$ for each player $i$; a transition function $P:\mathcal{S}\times\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{n}\to\Delta(\mathcal{S})$; reward functions $r_{i}:\mathcal{S}\times\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{n}\to\mathbb{R}$; and a common discount factor $\gamma\in[0,1)$. At each stage, all players simultaneously choose actions, receive individual rewards, and the game transitions to a new state.169169169Shapley (1953) introduced stochastic games in 1953 for the two-player zero-sum case, proving existence of the value via a contraction argument on the Bellman operator. The general-sum extension to $n$ players is due to Fink (1964) and Takahashi (1964).
@@ 9.1.2 Minimax-Q Learning
Minimax-Q converges to the minimax Q-values under Robbins-Monro learning rates ($\sum_{t}\alpha_{t}=\infty$, $\sum_{t}\alpha_{t}^{2}<\infty$) and infinite exploration of all state-action tuples.170170170The convergence proof extends the contraction argument for standard Q-learning. Because the zero-sum minimax operator is a contraction with modulus $\gamma$ under the $\ell^{\infty}$ norm, the stochastic approximation converges to the fixed point. See Littman (1994) and the general treatment in Szepesvári and Littman (1999). However, the algorithm sacrifices rationality: it plays the equilibrium strategy even against exploitable opponents.171171171In the soccer game of Littman (1994), minimax-Q won 53.7% against a hand-built opponent versus 26.1% for Q-learning. Against an adversarial challenger, Q-learning won 0% (its deterministic policy was fully predictable); minimax-Q won 37.5% through mixed strategies.
@@ 9.1.3 Nash-Q Learning
These conditions are restrictive. When stage games have multiple Nash equilibria, agents may select different equilibria for their backups, causing divergence.172172172In the experiments of Hu and Wellman (2003), a grid game with a unique equilibrium Q-function converged in 100% of trials under Nash-Q versus 20% under independent Q-learning; a game with three equilibrium Q-functions converged in only 68–90% of trials. Nash-Q requires each agent to observe all other agents’ rewards and to maintain Q-values over the joint action space $\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{n}$, with storage $O(n|\mathcal{S}|\prod_{i}|\mathcal{A}_{i}|)$, exponential in the number of players. Nash-Q reduces to standard Q-learning in the single-agent case.
@@ 9.1.4 The Convergence Problem
Two further algorithms deserve mention. Friend-or-Foe Q-learning (Littman, 2001) decomposes agents as cooperative (friend) or adversarial (foe), using $\max$ for friends and minimax for foes; it always converges but requires knowing the relationship type a priori.173173173Correlated-Q learning (Greenwald and Hall, 2003) generalizes both Nash-Q and Minimax-Q by using correlated equilibrium, a probability distribution over joint actions enforced by a correlating device. The set of correlated equilibria contains all Nash equilibria. Correlated-Q converges under conditions analogous to Nash-Q. The evolutionary perspective of Börgers and Sarin (1997) provides a deeper lens: reinforcement learning dynamics in matrix games converge to the replicator equation from evolutionary game theory. The cycling of Q-learning in games such as matching pennies is structurally identical to the cycling of replicator dynamics in Rock-Paper-Scissors games.
@@ 9.1.5 Simulation Study: Cournot and Bertrand Duopoly
Two canonical games from industrial organization serve as benchmarks. In Cournot duopoly, two firms choose quantities $q_{i}\in\{0,1,\ldots,9\}$ with inverse demand P(Q)=10QP(Q)=10-Q and marginal cost $c=1$; the unique Nash equilibrium is $q^{*}=3$ with profit $\pi^{*}=9$. In Bertrand duopoly with differentiated products, two firms choose prices $p_{i}\in\{0,1,\ldots,9\}$ with demand $d_{i}=10-2p_{i}+p_{j}$ and marginal cost $c=1$; the continuous Nash equilibrium is $p^{*}\approx 4.33$, which discretizes to $p^{*}=4$ with profit $\pi^{*}=18$. Both games have unique Nash equilibria in pure strategies.174174174Each configuration runs for 50,000 iterations across 20 seeds. Three algorithms compete: independent Q-learning (IQL), Nash-Q, and WoLF-PHC.
@@ 9.2 Counterfactual Regret Minimization
where $\pi^{\sigma}_{-i}(h)$ is the probability opponents reach $h$, and $\pi^{\sigma}(ha,z)$ is the probability of reaching terminal $z$ from $ha$. The counterfactual formulation weights by opponent reach $\pi^{\sigma}_{-i}(h)$ rather than joint reach $\pi^{\sigma}(h)$, ensuring non-zero updates even at rarely-visited information sets. Cumulative regret for action $a$ is $R^{T}(I,a)=\sum_{t=1}^{T}[v^{\sigma^{t}}(I,a)-v^{\sigma^{t}}(I)]$. CFR updates via regret matching:175175175Regret matching is an action selection rule where the probability of choosing action $a$ is proportional to the cumulative regret for not having chosen $a$ in the past (truncated at zero). Actions with high regret receive higher probability; actions with negative regret (having performed worse than average) receive zero probability.
@@
The average strategy $\bar{\sigma}^{T}$ converges to $\varepsilon$-Nash with $\varepsilon=O(|\mathcal{I}|\sqrt{|\mathcal{A}|}\Delta/\sqrt{T})$, where $\Delta$ is the range of payoffs (Zinkevich et al., 2008).176176176An $\varepsilon$-Nash equilibrium is a strategy profile where no player can improve her expected payoff by more than $\varepsilon$ through unilateral deviation. As $\varepsilon\to 0$, this converges to an exact Nash equilibrium.
CFR+ (Tammelin, 2014) replaces standard regret matching with Regret Matching+, which truncates cumulative regrets to zero after each update rather than only at action selection. The update becomes $R^{T+1}(I,a)=\max(R^{T}(I,a)+r^{T+1}(I,a),\,0)$, where $r^{T+1}(I,a)=v^{\sigma^{T+1}}(I,a)-v^{\sigma^{T+1}}(I)$ is the instantaneous counterfactual regret. CFR+ also weights iteration $t$ by $t$ when computing the average strategy. While vanilla CFR converges at $O(1/\sqrt{T})$, CFR+ empirically converges at O(1/T)O(1/T).177177177The $O(1/T)$ rate for CFR+ is empirically observed but not yet proven in full generality. Tammelin (2014) conjectured this rate based on extensive experiments across poker variants. CFR+ enabled Bowling et al. (2015) to essentially solve heads-up limit Texas hold’em, a game with $3.16\times 10^{17}$ states, the first non-trivial imperfect-information game played competitively by humans to be essentially solved. Their program Cepheus achieved exploitability below 1 mbb/g (milli-big-blind per game).
@@ 9.3.2 Neural Fictitious Self-Play
NFSP (Heinrich and Silver, 2016) combines fictitious play178178178Fictitious play (Brown, 1951) is a classical learning rule where each player best-responds to the empirical distribution of opponents’ past actions. Under certain conditions (e.g., two-player zero-sum games), the time-average strategies converge to Nash equilibrium. (Brown, 1951) with deep Q-learning.179179179Deep Q-Network (DQN) (Mnih et al., 2015) approximates the Q-function with a neural network, using experience replay and a target network to stabilize training. The target network $\theta^{-}$ in the loss function is a delayed copy of the main network, updated periodically. Each player maintains two networks: a best-response network $Q_{\theta}$ trained via DQN, and an average strategy network $\bar{\pi}_{\phi}$ trained via supervised learning on the best-response actions:
@@ 9.3.3 Poker Results
These methods achieved superhuman performance in poker. Deep CFR attained exploitability of 37 mbb/g180180180Exploitability measures how far a strategy is from Nash equilibrium, defined as the maximum expected gain an adversary could achieve by best-responding. Zero exploitability means the strategy is unexploitable (Nash). In poker, exploitability is measured in milli-big-blinds per game (mbb/g). in heads-up flop hold’em (Brown et al., 2019). Libratus defeated top human professionals in heads-up no-limit hold’em using nested subgame solving with CFR (Brown and Sandholm, 2018). Pluribus extended to six-player no-limit hold’em (Brown and Sandholm, 2019). The game tree of heads-up no-limit hold’em contains $\sim 10^{161}$ states; these results demonstrate that CFR-based methods scale to economically relevant game sizes.
@@ 9.4 The Coase Conjecture
Coase (1972) conjectured that a durable goods monopolist181181181A durable good provides utility over multiple periods (e.g., a car, appliance, or software license) rather than being consumed immediately. Unlike non-durable goods, durable goods create intertemporal competition, as the seller at time $t$ competes with her own future self at $t+1$. loses market power when buyers are patient. Unable to commit to future prices, the seller competes with her future self, eroding rents. As the inter-offer interval shrinks ($\delta\to 1$), price collapses to marginal cost and all surplus is extracted by buyers. Gul et al. (1986) formalized this for the gap case, where the seller’s cost is strictly below all buyer valuations, guaranteeing trade and a unique stationary equilibrium.182182182Stokey (1981) showed that a monopolist facing rational consumers cannot price discriminate intertemporally; Bulow (1982) proved that in the no-gap case, the monopolist prefers renting to selling. The no-gap case admits multiple equilibria with more complex dynamics.
@@ 9.4.2 Equilibrium Analysis
For $\pi<\pi^{*}$, the seller pools (offers $v_{L}$ immediately). For $\pi>\pi^{*}$, the seller screens (offers $P^{*}$, then $v_{L}$ if rejected).183183183In a screening equilibrium, the seller uses price to separate buyer types: high-value buyers accept immediately at a high price, while low-value buyers reject and receive a lower offer. In a pooling equilibrium, the seller offers a single price that all types accept. The gap case ($0=c<v_{L}$) guarantees trade in equilibrium. The Coase conjecture manifests as $\delta\to 1$, where $P^{*}(\delta)\to v_{L}$ and the seller cannot extract surplus from high types.

### 24 Mar 2026 v5

@@ 8.3.2 Fitted Policy Iteration for Combinatorial Auctions
The mechanism’s policy maps from state to a price vector over available items. The key technical contribution is making the auction clearing differentiable. In a standard auction, the allocation is an argmax over bids (non-differentiable), and payments depend discontinuously on the allocation. Ravindranath et al. (2024) replace the hard bundle selection with a softmax relaxation,168168168The softmax relaxation replaces the hard allocation $\mathop{\it argmax}_{i}b_{i}$ with a soft allocation $\exp(b_{i}/\tau)/\sum_{j}\exp(b_{j}/\tau)$, which is differentiable and approaches the hard allocation as $\tau\to 0$. enabling analytical gradient computation through the mechanism. The actor loss is the negative expected revenue, and gradients flow directly through the softmax-relaxed allocation and payment rules. The paper explicitly avoids REINFORCE-style estimators, noting that analytical gradients overcome the sample inefficiency and high variance of score-function methods.
@@
Experiments on settings with additive and combinatorial valuations show that the learned mechanisms achieve up to 13% higher revenue than item-wise Myerson optimal auctions, with the largest gains in combinatorial settings where bundle complementarities make item-wise pricing suboptimal. Standard RL baselines (PPO, SAC) are also outperformed, confirming the advantage of exploiting differentiable structure.169169169DDPG was also tested but found to be unstable. DQN is not applicable to continuous price-setting.
@@ 8.4 Macroeconomic Models
Reinforcement learning and deep learning are increasingly used to solve high-dimensional macroeconomic models where grid-based dynamic programming is infeasible. Heterogeneous agent models, in which a distribution of agents with different wealth levels, productivities, or beliefs interact through markets, generate state spaces that grow with the number of agent types and asset positions. Maliar et al. (2021) demonstrate that deep neural networks can approximate policy and value functions in dynamic economic models, achieving accuracy comparable to established projection methods while scaling to problems with dozens of state variables. Fernández-Villaverde et al. (2023) solve a model with financial frictions and an endogenous wealth distribution using deep learning, obtaining global solutions to a problem where perturbation methods fail due to strong nonlinearities. Fernández-Villaverde et al. (2024) provide a systematic treatment of deep learning methods for high-dimensional dynamic programming problems in economics, covering both single-agent and equilibrium settings. Atashbar and Shi (2023) apply deep deterministic policy gradient (DDPG)170170170DDPG extends Q-learning to continuous action spaces by learning a deterministic policy network alongside a Q-function critic. to a real business cycle model, demonstrating that model-free RL can recover near-optimal consumption and investment policies without deriving optimality conditions such as the Euler equation. Moll (2025) argues that rational expectations equilibria in heterogeneous agent models are computationally intractable and proposes reinforcement learning as a more tractable alternative for modeling how agents form beliefs and make decisions; see also the textbook treatment in Zhao (2025). This is a rapidly growing area; readers seeking a comprehensive methodological treatment are directed to the cited papers and the references therein.
@@ 9.1.1 The Stochastic Game Framework
An $n$-player stochastic game $\Gamma=(n,\mathcal{S},\mathcal{A}_{1},\ldots,\mathcal{A}_{n},P,r_{1},\ldots,r_{n},\gamma)$ consists of a finite state space $\mathcal{S}$; finite action sets $\mathcal{A}_{i}$ for each player $i$; a transition function $P:\mathcal{S}\times\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{n}\to\Delta(\mathcal{S})$; reward functions $r_{i}:\mathcal{S}\times\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{n}\to\mathbb{R}$; and a common discount factor $\gamma\in[0,1)$. At each stage, all players simultaneously choose actions, receive individual rewards, and the game transitions to a new state.171171171Shapley (1953) introduced stochastic games in 1953 for the two-player zero-sum case, proving existence of the value via a contraction argument on the Bellman operator. The general-sum extension to $n$ players is due to Fink (1964) and Takahashi (1964).
@@ 9.1.2 Minimax-Q Learning
Minimax-Q converges to the minimax Q-values under Robbins-Monro learning rates ($\sum_{t}\alpha_{t}=\infty$, $\sum_{t}\alpha_{t}^{2}<\infty$) and infinite exploration of all state-action tuples.172172172The convergence proof extends the contraction argument for standard Q-learning. Because the zero-sum minimax operator is a contraction with modulus $\gamma$ under the $\ell^{\infty}$ norm, the stochastic approximation converges to the fixed point. See Littman (1994) and the general treatment in Szepesvári and Littman (1999). However, the algorithm sacrifices rationality: it plays the equilibrium strategy even against exploitable opponents.173173173In the soccer game of Littman (1994), minimax-Q won 53.7% against a hand-built opponent versus 26.1% for Q-learning. Against an adversarial challenger, Q-learning won 0% (its deterministic policy was fully predictable); minimax-Q won 37.5% through mixed strategies.
@@ 9.1.3 Nash-Q Learning
These conditions are restrictive. When stage games have multiple Nash equilibria, agents may select different equilibria for their backups, causing divergence.174174174In the experiments of Hu and Wellman (2003), a grid game with a unique equilibrium Q-function converged in 100% of trials under Nash-Q versus 20% under independent Q-learning; a game with three equilibrium Q-functions converged in only 68–90% of trials. Nash-Q requires each agent to observe all other agents’ rewards and to maintain Q-values over the joint action space $\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{n}$, with storage $O(n|\mathcal{S}|\prod_{i}|\mathcal{A}_{i}|)$, exponential in the number of players. Nash-Q reduces to standard Q-learning in the single-agent case.
@@ 9.1.4 The Convergence Problem
Two further algorithms deserve mention. Friend-or-Foe Q-learning (Littman, 2001) decomposes agents as cooperative (friend) or adversarial (foe), using $\max$ for friends and minimax for foes; it always converges but requires knowing the relationship type a priori.175175175Correlated-Q learning (Greenwald and Hall, 2003) generalizes both Nash-Q and Minimax-Q by using correlated equilibrium, a probability distribution over joint actions enforced by a correlating device. The set of correlated equilibria contains all Nash equilibria. Correlated-Q converges under conditions analogous to Nash-Q. The evolutionary perspective of Börgers and Sarin (1997) provides a deeper lens: reinforcement learning dynamics in matrix games converge to the replicator equation from evolutionary game theory. The cycling of Q-learning in games such as matching pennies is structurally identical to the cycling of replicator dynamics in Rock-Paper-Scissors games.
@@ 9.1.5 Simulation Study: Cournot and Bertrand Duopoly
Two canonical games from industrial organization serve as benchmarks. In Cournot duopoly, two firms choose quantities $q_{i}\in\{0,1,\ldots,9\}$ with inverse demand P(Q)=10QP(Q)=10-Q and marginal cost $c=1$; the unique Nash equilibrium is $q^{*}=3$ with profit $\pi^{*}=9$. In Bertrand duopoly with differentiated products, two firms choose prices $p_{i}\in\{0,1,\ldots,9\}$ with demand $d_{i}=10-2p_{i}+p_{j}$ and marginal cost $c=1$; the continuous Nash equilibrium is $p^{*}\approx 4.33$, which discretizes to $p^{*}=4$ with profit $\pi^{*}=18$. Both games have unique Nash equilibria in pure strategies.176176176Each configuration runs for 50,000 iterations across 20 seeds. Three algorithms compete: independent Q-learning (IQL), Nash-Q, and WoLF-PHC.
@@ 9.2 Counterfactual Regret Minimization
where $\pi^{\sigma}_{-i}(h)$ is the probability opponents reach $h$, and $\pi^{\sigma}(ha,z)$ is the probability of reaching terminal $z$ from $ha$. The counterfactual formulation weights by opponent reach $\pi^{\sigma}_{-i}(h)$ rather than joint reach $\pi^{\sigma}(h)$, ensuring non-zero updates even at rarely-visited information sets. Cumulative regret for action $a$ is $R^{T}(I,a)=\sum_{t=1}^{T}[v^{\sigma^{t}}(I,a)-v^{\sigma^{t}}(I)]$. CFR updates via regret matching:177177177Regret matching is an action selection rule where the probability of choosing action $a$ is proportional to the cumulative regret for not having chosen $a$ in the past (truncated at zero). Actions with high regret receive higher probability; actions with negative regret (having performed worse than average) receive zero probability.
@@
The average strategy $\bar{\sigma}^{T}$ converges to $\varepsilon$-Nash with $\varepsilon=O(|\mathcal{I}|\sqrt{|\mathcal{A}|}\Delta/\sqrt{T})$, where $\Delta$ is the range of payoffs (Zinkevich et al., 2008).178178178An $\varepsilon$-Nash equilibrium is a strategy profile where no player can improve her expected payoff by more than $\varepsilon$ through unilateral deviation. As $\varepsilon\to 0$, this converges to an exact Nash equilibrium.
CFR+ (Tammelin, 2014) replaces standard regret matching with Regret Matching+, which truncates cumulative regrets to zero after each update rather than only at action selection. The update becomes $R^{T+1}(I,a)=\max(R^{T}(I,a)+r^{T+1}(I,a),\,0)$, where $r^{T+1}(I,a)=v^{\sigma^{T+1}}(I,a)-v^{\sigma^{T+1}}(I)$ is the instantaneous counterfactual regret. CFR+ also weights iteration $t$ by $t$ when computing the average strategy. While vanilla CFR converges at $O(1/\sqrt{T})$, CFR+ empirically converges at O(1/T)O(1/T).179179179The $O(1/T)$ rate for CFR+ is empirically observed but not yet proven in full generality. Tammelin (2014) conjectured this rate based on extensive experiments across poker variants. CFR+ enabled Bowling et al. (2015) to essentially solve heads-up limit Texas hold’em, a game with $3.16\times 10^{17}$ states, the first non-trivial imperfect-information game played competitively by humans to be essentially solved. Their program Cepheus achieved exploitability below 1 mbb/g (milli-big-blind per game).
@@ 9.3.2 Neural Fictitious Self-Play
NFSP (Heinrich and Silver, 2016) combines fictitious play180180180Fictitious play (Brown, 1951) is a classical learning rule where each player best-responds to the empirical distribution of opponents’ past actions. Under certain conditions (e.g., two-player zero-sum games), the time-average strategies converge to Nash equilibrium. (Brown, 1951) with deep Q-learning.181181181Deep Q-Network (DQN) (Mnih et al., 2015) approximates the Q-function with a neural network, using experience replay and a target network to stabilize training. The target network $\theta^{-}$ in the loss function is a delayed copy of the main network, updated periodically. Each player maintains two networks: a best-response network $Q_{\theta}$ trained via DQN, and an average strategy network $\bar{\pi}_{\phi}$ trained via supervised learning on the best-response actions:
@@ 9.3.3 Poker Results
These methods achieved superhuman performance in poker. Deep CFR attained exploitability of 37 mbb/g182182182Exploitability measures how far a strategy is from Nash equilibrium, defined as the maximum expected gain an adversary could achieve by best-responding. Zero exploitability means the strategy is unexploitable (Nash). In poker, exploitability is measured in milli-big-blinds per game (mbb/g). in heads-up flop hold’em (Brown et al., 2019). Libratus defeated top human professionals in heads-up no-limit hold’em using nested subgame solving with CFR (Brown and Sandholm, 2018). Pluribus extended to six-player no-limit hold’em (Brown and Sandholm, 2019). The game tree of heads-up no-limit hold’em contains $\sim 10^{161}$ states; these results demonstrate that CFR-based methods scale to practically relevant game sizes.
@@ 9.4 The Coase Conjecture
Coase (1972) conjectured that a durable goods monopolist183183183A durable good provides utility over multiple periods (e.g., a car, appliance, or software license) rather than being consumed immediately. Unlike non-durable goods, durable goods create intertemporal competition, as the seller at time $t$ competes with her own future self at $t+1$. loses market power when buyers are patient. Unable to commit to future prices, the seller competes with her future self, eroding rents. As the inter-offer interval shrinks ($\delta\to 1$), price collapses to marginal cost and all surplus is extracted by buyers. Gul et al. (1986) formalized this for the gap case, where the seller’s cost is strictly below all buyer valuations, guaranteeing trade and a unique stationary equilibrium.184184184Stokey (1981) showed that a monopolist facing rational consumers cannot price discriminate intertemporally; Bulow (1982) proved that in the no-gap case, the monopolist prefers renting to selling. The no-gap case admits multiple equilibria with more complex dynamics.
@@ 9.4.2 Equilibrium Analysis
For $\pi<\pi^{*}$, the seller pools (offers $v_{L}$ immediately). For $\pi>\pi^{*}$, the seller screens (offers $P^{*}$, then $v_{L}$ if rejected).185185185In a screening equilibrium, the seller uses price to separate buyer types: high-value buyers accept immediately at a high price, while low-value buyers reject and receive a lower offer. In a pooling equilibrium, the seller offers a single price that all types accept. The gap case ($0=c<v_{L}$) guarantees trade in equilibrium. The Coase conjecture manifests as $\delta\to 1$, where $P^{*}(\delta)\to v_{L}$ and the seller cannot extract surplus from high types.